# eprint Nisho-2-2025

### v4

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B_{0}$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the power $P$ is estimated as $P\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. The signal of the axion can be detected by observing electric current in parallel LC circuit with quality factor $Q\geq 2$. The detection of dark matter axions in our method may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 2020mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
On the other hand, the two suppression factors do not arise in materials with small conductivity such as semiconductor, Especially, for the cylinder sample with small conductivity $\sigma=10^{-3}$eV, the induced current mainly flows the bulk, and the resulting electric current $I$ and power $P$ become large. It would be detectable at low temperatures ( $T\sim 10\text{mK}$ ). We show that when the axion mass $m_{a}=10^{-4}$eV,
$I(\sigma=10^{-3}\text{eV})$ $\simeq$ $2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (1)
$P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
@@
$P(\sigma=10^{4}\text{eV})\simeq 5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (3)
@@
In the observation, the electric $I(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ where the resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder and $\delta\omega=10^{-6}m_{a}$,
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 4.2\times 10^{-4}g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\mbox{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)\Big(\frac{B}{15\text{T}}\Big)\Big(\frac{10^{-4}\text{eV}}{m_{a}}\Big)^{1/2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (4)
The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=100$seconds. In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.6\times 10^{3}$, yielding $I(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 6.7\times 10^{-1}$.
The detection of the current is observed using parallel LC circuit in which the current is amplified with resonance by tuning parameters of the circuit. Even when we have small enhancement factor $Q=10$, the signal-noise ratio becomes larger than $1$; $QI(\sigma=10^{-3}\text{eV})\sqrt{\delta\omega\delta t_{ob}/2\pi}/I_{n}\simeq 4.6$. Consequently, the signal of the axion can be detected by our method.
Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of the so-called dark photon. The electromagnetic coupling of the dark photon produces an effect similar to that of the axion.
@@
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)=1/g_{a\gamma\gamma}$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with the depth $\delta$ of the cylinder. It is coincident with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
@@
$P$ takes the maximal value at $y=\sigma/m_{a}=\epsilon$ and $P(\epsilon,y=\epsilon)\propto\epsilon^{-1}$. Thus, it is favorable to take the conductivity $\sigma=\epsilon m_{a}$. Because we do not know the value of the axion mass $m_{a}$, the value $\sigma=\epsilon m_{a}$ is unknown. But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$. In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV with $\epsilon\sim 10$. Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection. Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
@@
$P(\sigma=10^{-3}\text{eV})\simeq 5.7\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (24)
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 30$ ( R=6R=6cm ). Obviously, $P(\sigma=10^{-3}\text{eV})$ in eq.(24) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$ (27)
where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(16).
The point causing the large current $I(\sigma=10^{-3}\text{eV})$ is that electric field $E$ inside of the cylinder with small conductivity, e.g. $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV. Additionally, the current flows in the whole of the cylinder, not just in the surface. The facts make the large electric current $I(\sigma=10^{-3}\text{eV})\sim 10^{-14}$A.
@@
$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-4}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (30)
@@
$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\mbox{eV}}\Big)^{1/2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (31)
These powers can be detected using a parallel LC circuit, in which electric current in coil is measured. By tuning the parameters of the LC circuit, the electric current at resonance is enhanced by a factor $Q=R_{c}/m_{a}L_{c}$ where $L_{c}$ denotes inductance of the inductor and resistance $R_{c}=R$ of the cylinder. The cylinder forms the resistor of the circuit. The resonance condition is given by $m_{a}=1/\sqrt{C_{c}L_{c}}$ where $C_{c}$ is capacitance of the capacitor.
In order to observe the electric current $I(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-14}$A, we compare it with the thermal noise $I_{n}$,
@@
with the resistance $R_{c}=L/(\sigma\pi R^{2})\simeq 1.44(L/100\text{cm})(6\text{cm}/R^{2})\simeq 540\Omega$. Then, a signal-to-noise ratio at resonance is given by
@@
$\frac{QI(\sigma=10^{-3}\text{eV})}{I_{n}({\sigma=10^{-3}\text{eV})}}$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq$ (33)

### v5

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B_{0}$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the current $I$ is estimated as $I(\sigma=10^{-3}\text{eV})\simeq 2.8\times 10^{-14}\text{A}g_{\gamma}\big(R/6\text{cm}\big)^{2}\big(\sigma/10^{-3}\text{eV}\big)\big(B_{0}/15\text{T}\big)\big(10/\epsilon\big)\big(\rho_{a}/0.3\rm GeVcm^{-3}\big)^{1/2}$ for $m_{a}=10^{-4}$eV, with radius $R$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. Because the current is proportional to $R^{2}$, using large sample with R=80R=80cm, we have large signal-noise ratio ( >1>1 ) even in temperature T=4T=4K, $I(\sigma=10^{-3}\text{eV})/I_{n}({\sigma=10^{-3}\text{eV})}\times\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 1.1g_{\gamma}(4\text{K}/T)^{1/2}(L/100\text{cm})^{1/2}(R/80\text{cm})(B_{0}/7\mbox{T})(\rho_{a}/0.3\rm GeVcm^{-3})^{1/2}(\delta t_{ob}/10^{3}\,\text{s})^{1/2}$ for $m_{a}=10^{-4}\text{eV}$ with $\epsilon=10$ and $\sigma=\epsilon m_{a}$, where thermal noise is $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with $\delta\omega=10^{-6}m_{a}$ and resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder with length $L$. Although a superconducting solenoid sufficiently large to accommodate such a sample is required, the detection of dark matter axions in our proposal may be feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature $\sim 1$K. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
On the other hand, the two suppression factors do not arise in materials with small conductivity such as semiconductor. One is that we have no suppression factor, $\sqrt{m_{a}/\sigma}$ for the cylinder sample with small conductivity $\sigma=10^{-3}$eV. The other one is that the induced current mainly flows the bulk. The presence of the bulk current is a distinctive feature of electromagnetic fields coupled with axion. The resulting electric current $I$ and power $P$ become large with large sample. It would be detectable at low temperatures ( $T\sim 1\text{K}$ ).
We show that when the axion mass $m_{a}=10^{-4}$eV,
$I(\sigma=10^{-3}\text{eV})$ $\simeq$ $2.9\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}$ (1)
$P(\sigma=10^{-3}\text{eV})$ $\simeq$ $5.3\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
@@
$P(\sigma=10^{4}\text{eV})\simeq 5.4\times 10^{-31}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)\sqrt{\frac{m_{a}}{10^{-4}\mbox{eV}}}\sqrt{\frac{10^{4}\mbox{eV}}{\sigma}}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (3)
@@
In the observation, the electric current $I(\sigma=10^{-3}\text{eV})$ should be compared with thermal noise $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ where the resistance $R_{c}=L/(\sigma\pi R^{2})$ of the cylinder and $\delta\omega=10^{-6}m_{a}$,
$\frac{I\big(\sigma=(10^{-3}\text{eV}/\sqrt{3}\big))}{I_{n}(\sigma=\big(10^{-3}\text{eV}/\sqrt{3})\big)}$ $\simeq$ $5.0\times 10^{-4}g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\mbox{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)$ (4)
$\times$ $\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{3/2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$
with $m_{a}=10^{-4}$eV, where we have taken $\sigma=\epsilon m_{a}/\sqrt{3}=(10^{-3}\text{eV}/\sqrt{3})$, which maximizes the ratio $I/I_{n}$.
The noise is further reduced by considering a frequency width $\delta\omega=10^{-6}m_{a}$ and an observation time $\delta t_{ob}=10^{3}$seconds. In this case, the noise decreases by the factor $\sqrt{\delta\omega\delta t_{ob}/2\pi}\simeq 4.9\times 10^{3}$, yielding
$\frac{I\big(\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}}{I_{n}\big(\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)}\simeq 2.5.$ (5)
Achieving an electrical conductivity of order $10^{-3}$eV in semiconductors at 2020mK is challenging, but becomes feasible at 44K. Although thermal noise is higher at this temperature, the signal can be correspondingly enhanced when we take large radius $R$. Since the induced current scales with the square of the sample radius, the use of larger samples, i.e. R=50R=50cm, significantly increases the signal. Consequently, a large signal-to-noise ratio can be achieved even at 44K.
$\frac{I\big(\sigma=\big(10^{-3}\text{eV}/\sqrt{3})\big)}{I_{n}\big({\sigma=(10^{-3}\text{eV}/\sqrt{3})\big)}}$ $\times$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}\simeq 0.98g_{\gamma}\Big(\frac{4\text{K}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{50\text{cm}}\Big)$ (6)
$\times$ $\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sqrt{3}\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}.$
Our method, which employs a cylinder with low electrical conductivity, is also applicable to the detection of axion like particle ( ALP ) or so-called dark photon. The electromagnetic coupling of theses particles produces a signal similar to that of the axion.
In the next section, we briefly explain the property of QDC axion coupled with electromagnetic fields. Using the coupling, we solve electromagnetic fields generated inside the cylinder under strong magnetic field and calculate electric power $P$ and current $I$ in section (III). In the section (IV), we discuss realistic choices of materials forming the large cylinder R=80R=80cm and temperature 44K cooling the cylinder. We present a figure in which detectable regions of ALPs are shown in $g_{a\gamma\gamma}-m_{a}$ plane. We also show that ALPs with extremely small mass, e.g. $10^{-14}$eV can be detectable with the choice of electric conductivity being of the order of the mass.
@@
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)=1/g_{a\gamma\gamma}$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with the depth $\delta$ of the cylinder. It is coincident with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$ and that $b(y)\to\sqrt{\sigma/m_{a}}\exp(i\pi/4)$ in the limit $y=\sigma/m_{a}\gg 1$.
@@
$P$ takes the maximal value at $y=\sigma/m_{a}=\epsilon$ and $P(\epsilon,y=\epsilon)\propto\epsilon^{-1}$. Thus, it is favorable to take the conductivity $\sigma=\epsilon m_{a}$. Because we do not know the value of the axion mass $m_{a}$, the value $\sigma=\epsilon m_{a}$ is unknown. But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$. Otherwise, the dependence of $P$ on $m_{a}$ or $\sigma$ is not simply scaling; $P\propto\sigma m_{a}^{2}/(\epsilon^{2}m_{a}^{2}+\sigma^{2})$. In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, it is sufficient for the axion detection to take the value $\sigma\sim 1.0\times 10^{-3}$eV with $\epsilon\sim 10$. Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection. Hereafter, we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ to estimate the power $P$. Later we find that $\sigma=\epsilon m_{a}/\sqrt{3}$ maximizes the ratio of electric current $I$ induced by axion to thermal noise $I_{n}$. Please refer Fig.2 in which the dependence of $P(\sigma,m_{a})$ and the current $I(\sigma,m_{a})$ on $\sigma$ and $m_{a}$ is shown.
@@
$P(\sigma=10^{-3}\text{eV})\simeq 5.3\times 10^{-27}\mbox{W}g_{\gamma}^{2}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{6\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (26)
with $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 30$ ( $x=m_{a}R=10^{-3}\text{eV}\times 6\text{cm}\simeq 30$ ). Obviously, $P(\sigma=10^{-3}\text{eV})$ in eq.(26) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma\epsilon|E_{a}|S}{(\epsilon^{2}+y^{2})}\simeq 2.9\times 10^{-14}\text{A}g_{\gamma}\Big(\frac{R}{6\text{cm}}\Big)^{2}\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{B_{0}}{15\text{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2},$ (29)
where the electric field $E=Re(E^{\prime})=\epsilon|E_{a}|/(\epsilon^{2}+y^{2})=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(18).
The point causing the large current $I(\sigma=10^{-3}\text{eV})$ is that electric field $E$ inside of the cylinder with small conductivity, e.g. $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV. Additionally, the current flows in the bulk of the cylinder, not just in the surface. The facts make the large electric current $I(\sigma=10^{-3}\text{eV})\sim 10^{-14}$A.
Figure 2 shows the dependence of the power $P(\sigma,m_{a})$ and current $I(\sigma,m_{a})$ on the axion mass $m_{a}$ and electrical conductivity $\sigma$ when $\epsilon=10$. The vertical axis is given in arbitrary units. It is remarkable feature that as the axion mass $m_{a}$ increases, increasing the electrical conductivity $\sigma$ correspondingly leads to larger values of both $P(\sigma,m_{a})$ and $I(\sigma,m_{a})$.
Figure 2: Dependence of power $P(\sigma,m_{a})$ and current $I(\sigma,m_{a})$ on $\sigma$ and $m_{a}$ when $\epsilon=10$.
@@
$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-3}\text{eV}}\big)\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{3/2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (32)
@@
$\Big(\frac{B_{0}}{15\mbox{T}}\Big)^{2}\Big(\frac{2\big(\frac{\sigma}{10^{-2}\text{eV}}\big)\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{3/2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-3}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-2}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\Big(\frac{\delta t_{ob}}{100\,\text{s}}\Big)^{1/2}.$ (33)
SN ratio of the power is less than $10^{-2}$, but we can show that SN ratio of electric current $I$ is bigger than $1$.
We should make a comment that we have taken $y\equiv\sigma/m_{a}=\epsilon$ in order to maximize the SN ratio $P/P_{n}$. On the other hand, to maximize SN ratio $I/I_{n}$ of the electric current to thermal noise $I_{n}$, we need to take $y=\epsilon/\sqrt{3}$, because $I/I_{n}\propto yy^{-1/2}/(\epsilon^{2}+y^{2})$: $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with the resistance of the cylinder, $R_{c}=L/(\sigma\pi R^{2})$. However, we take $y=\epsilon$ for simplicity in the subsequent calculations. It simply causes slightly smaller value of the ratio $I/I_{n}$; $y^{1/2}/(\epsilon^{2}+y^{2})\simeq 1.58\times 10^{-3}$ with $y=\epsilon=10$, while $y^{1/2}/(\epsilon^{2}+y^{2})\simeq 1.8\times 10^{-3}$ with $y=\epsilon/\sqrt{3}$.
Because thermal noise $I_{n}$ is
@@
with the resistance $R_{c}\simeq 1.44(L/100\text{cm})(6\text{cm}/R)^{2}(10^{-3}\text{eV}/\sigma)\simeq 540\Omega$, the SN ratio is
@@
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}({\sigma=10^{-3}\text{eV})}}$ $\sqrt{\frac{\delta\omega\delta t_{ob}}{2\pi}}$ $\simeq 2.1g_{\gamma}\Big(\frac{20\text{mK}}{T}\Big)^{1/2}\Big(\frac{L}{100\text{cm}}\Big)^{1/2}\Big(\frac{R}{6\text{cm}}\Big)$ (35)
$\times\Big(\frac{B_{0}}{15\text{T}}\Big)$ $\Big(\frac{2\big(\frac{\epsilon}{10}\big)\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{1/2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}}{\big(\frac{\epsilon}{10}\big)^{2}\big(\frac{m_{a}}{10^{-4}\text{eV}}\big)^{2}+\big(\frac{\sigma}{10^{-3}\text{eV}}\big)^{2}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)^{1/2}\Big(\frac{\delta t_{ob}}{1000\,\text{s}}\Big)^{1/2}$
It is the ratio corresponding to axion mass $m_{a}=10^{-4}$eV, while it is given by